Monitors manufactured by TSI Electronics have life spans that have a normal distribution with a variance of 1,690,000 and a mean life span of 17,000 hours. If a monitor is selected at random, find the probability that the life span of the monitor will be more than 14,919 hours. Round your answer to four decimal places.
Answer #1

Let, XNNIW=19,000 ,5*1,690,000) Now, plX314,919)= pixell 114,919-19,000 1/1,690,000 =P(Z-1.60] [0911-5 % 10-13 (09.1-70-15 20